Gross Corporation adopted the dollar-value LIFO method of inventory valuation on December 31, 2016. Its inventory at that datGross Corporation adopted the dollar-value LIFO method of inventory valuation on December 31, 2016. Its inventory at that date was $1100000 and the relevant price index was 100. Information regarding inventory for subsequent years is as follows: Date Inventory at Current Prices Current Price Index December 31, 2017 $1284000 107 December 31, 2018 1450000 125 December 31, 2019 1625000 130 What is the cost of the ending inventory at December 31, 2019 under dollar-value LIFO?

*PLEASE EXPLAIN IN DETAIL, LIKE WHERE THE NUMBERS COME FROM*
Answer #1
Ending inventory
2016 1100000
2017 1284000/1.07 = 1200000 (1100000+100000*1.07) = 1207000
2018 1450000/1.25 = 1160000 (1100000+60000*1.07) = 1164200
2019 1625000/1.30 = 1250000 (1164200+90000*1.30) = 1281200

So answer is a) $1281200

Note : Please rating